# One Patch to Caption Them All: A Unified Zero-Shot Captioning Framework

### v2

@@ 9 Modality Gap: Projection to Textual Space vs Training with Noise
Various works Gu et al. (2023); Nukrai et al. (2022) proposed zero-shot image captioning solutions based on noise injection during the training of the text decoder. Through this strategy, the trained decoders are more effective in understanding semantic representations, even when those are not coming from the text modality. To implement this strategy in our framework, we trained the textual decoder on the same collection of captions as for the memory bank-based configuration. We adopted Talk2DINO Barsellotti et al. (2024) textual space for the decoder input space, which is aligned to DINOv2 Oquab et al. (2023) with registers Jose et al. (2024). Following the setting of Gu et al. (2023), we added Gaussian noise with $\sigma^{2}=0.08$ to the textual embeddings while leaving the other parameters unchanged (as defined in §6). In the next paragraphs, we report and compare the results for each task of Talk2DINO within our framework with the memory bank ($\mathcal{M}\textit{emory}$) and with the training with noise ($\mathcal{N}$oise).
@@
In Table 7, we compare the two modality gap mitigation strategies across multiple captioning tasks, and also report the performance of a baseline without any mitigation (no mitig.). The baseline consistently underperforms compared to both the $\mathcal{M}\textit{emory}$ and $\mathcal{N}$oise configurations, indicating that, like other contrastively learned image-text encoders Liang et al. (2022), Talk2DINO is also affected by the modality gap. These results highlight the importance of explicitly addressing this gap to achieve strong captioning performance. For Trace Captioning, the $\mathcal{M}\textit{emory}$ method is slightly more effective in the semantic metric RefPAC-S, while the $\mathcal{N}$oise variant achieves marginally better scores in CIDEr, ROUGE-L, METEOR, and BLEU@4, with a minimal gap between the two approaches. In Dense Captioning, the $\mathcal{M}\textit{emory}$ model consistently outperforms the $\mathcal{N}$oise model across all metrics. Similarly, for Region-Set Captioning, both methods achieve strong results, but the $\mathcal{M}\textit{emory}$ method shows a clearer advantage, particularly in tasks closer to the patch level. Finally, in Image Captioning, the performance gap between the two architectures narrows, especially on the Flickr30k test split. In this scenario, the $\mathcal{M}\textit{emory}$ method performs significantly better when applied to the CLS token, whereas patch aggregation produces comparable results. However, the metrics reveal conflicting trends across different datasets.
@@ 10 Trace Captioning Benchmark Generation
We construct our Trace Captioning dataset from the Localized Narratives dataset Pont-Tuset et al. (2020a). This dataset consists of mouse traces and their corresponding speech transcriptions, where annotators describe objects in images while moving the mouse pointer over them.
@@
      {internallinenumbers*} I have image descriptions derived from spoken narratives. These need to be rewritten as concise, stand-alone captions in the style of the image-caption datasets. Follow these rules:      {internallinenumbers*} - Remove unnecessary narrative phrases like "we can see," "there is," "in this image," etc.    - Ensure the caption is standalone and descriptive.    - Use simple, objective language that highlights key elements.    - Keep it concise—just a single phrase.    - Follow the classical style of caption datasets.    {internallinenumbers*} - If the description is vague, subjective, or does not describe a concrete visual element (e.g., "The image is taken indoor," "This image is blurred"), return `<INVALID>`.    - Wrap the output in `{}` and add nothing else.      ### **Examples:**    {internallinenumbers*} - **Input:** "We can see a young elephant stands which is near the water in a wooded area."    **Output:** {A young elephant stands near the water in a wooded area.}      {internallinenumbers*} - **Input:** "In this image I can see some young children kicking a soccer ball in a field."    {internallinenumbers*} **Output:** {A group of young children kicking a soccer ball around a field.}      {internallinenumbers*} - **Input:** "In the left of the image, we see a pole that has two green street signs on it."    **Output:** {A pole has two green street signs on it.}      {internallinenumbers*} - **Input:** "We can see two surfboards which are stuck in the sand along the seashore."    **Output:** {Two surfboards stuck in the sand along the seashore.}      {internallinenumbers*} - **Input:** "This image consists of a man which rides a wakeboard behind a boat."    **Output:** {A man rides a wakeboard behind a boat.}      {internallinenumbers*} - **Input:** "In the background, there are a bunch of sticky notes and a pair of scissors."    **Output:** {A bunch of sticky notes and a pair of scissors.}      {internallinenumbers*} - **Input:** "It looks like a sepia-toned photograph of a motorcycle underneath the shadow of a    tree."    {internallinenumbers*} **Output:** {A sepia-toned photograph of a motorcycle underneath the shadow of a tree.}      - **Input:** "There is a sky"    **Output:** {A sky.}      - **Input:** "She is smiling."    **Output:** {A smiling girl.}      - **Input:** "The image is taken indoor."    **Output:** {<INVALID>}      - **Input:** "This image is edited."    **Output:** {<INVALID>}      - **Input:** "The image is blurred."    **Output:** {<INVALID>}      - **Input:** "I think he is about to jump."    **Output:** {<INVALID>}      {internallinenumbers*} Now, rewrite the following captions accordingly. Wrap each in `{}` and add nothing else:    <INPUT CAPTION>
@@
Figure 5: Narrative vs. Trace Samples. The first column displays sample images from the Localized Narrative dataset Pont-Tuset et al. (2020b). The remaining three columns show the corresponding mouse traces, along with the captions generated by the LLM. Captions marked with <INVALID> are removed from the dataset.
@@ 11 More Qualitative Results
Additional qualitative results are shown in Figures 6 and 7. Note that the first rows of Figures 6 and 7 contain also qualitative results for single patch captioning, for which we do not have annotated data to report quantitative results.
As can be noticed in Figures 6 and 7, the Region-Set Captioning task tends to align more closely with image-level captioning rather than strictly focusing on localized regions. This is expected since the ground-truth captions in the COCO Entities dataset originate from the image-level annotations of COCO, as stated in Cornia et al. (2019).
@@
Figure 6: Qualitative results. We report four predictions of our model and compare baselines from the finer (top) to the coarser (bottom) task. For trace captioning examples, the trace time is color-coded from start (red) to end (yellow). DeCap = DeCap applied on the whole image. DeCap (Crop) = DeCap applied on cropped box. ZeroCap = ZeroCap Tewel et al. (2022) applied to the whole image. CLOSE = CLOSE Gu et al. (2023) applied to the whole image. Ours (CLIP + Mem.) = Our patch-based framework using CLIP as backbone and the projection as modality gap mitigation strategy. Ours (Talk2DINO + Mem.) = Our patch-based framework using Talk2DINO as backbone and the projection as modality gap mitigation strategy. GT = ground-truth caption.
@@
Figure 7: Qualitative results. We report four predictions of our model and compare baselines from the finer (top) to the coarser (bottom) task. For trace captioning examples, the trace time is color-coded from start (red) to end (yellow). DeCap = DeCap applied on the whole image. DeCap (Crop) = DeCap applied on cropped box. ZeroCap = ZeroCap Tewel et al. (2022) applied to the whole image. CLOSE = CLOSE Gu et al. (2023) applied to the whole image. Ours (CLIP + Mem.) = Our patch-based framework using CLIP as backbone and the projection as modality gap mitigation strategy. Ours (Talk2DINO + Mem.) = Our patch-based framework using Talk2DINO as backbone and the projection as modality gap mitigation strategy. GT = ground-truth caption.

### v3

@@ 12 Modality Gap: Projection to Textual Space vs Training with Noise
Various works [18, 43] proposed zero-shot image captioning solutions based on noise injection during the training of the text decoder. Through this strategy, the trained decoders are more effective in understanding semantic representations, even when those are not coming from the text modality. To implement this strategy in our framework, we trained the textual decoder on the same collection of captions as for the memory bank-based configuration. We adopted Talk2DINO [5] textual space for the decoder input space, which is aligned to DINOv2 [44] with registers [24]. Following the setting of [18], we added Gaussian noise with $\sigma^{2}=0.08$ to the textual embeddings while leaving the other parameters unchanged (as defined in §6). In the next paragraphs, we report and compare the results for each task of Talk2DINO within our framework with the memory bank ($\mathcal{M}\textit{emory}$) and with the training with noise ($\mathcal{N}$oise).
@@
In Table 10, we compare the two modality gap mitigation strategies across multiple captioning tasks, and also report the performance of a baseline without any mitigation (no mitig.). The baseline consistently underperforms compared to both the $\mathcal{M}\textit{emory}$ and $\mathcal{N}$oise configurations, indicating that, like other contrastively learned image-text encoders [36], Talk2DINO is also affected by the modality gap. These results highlight the importance of explicitly addressing this gap to achieve strong captioning performance. For Trace Captioning, the $\mathcal{M}\textit{emory}$ method is slightly more effective in the semantic metric RefPAC-S, while the $\mathcal{N}$oise variant achieves marginally better scores in CIDEr, ROUGE-L, METEOR, and BLEU@4, with a minimal gap between the two approaches. In Dense Captioning, the $\mathcal{M}\textit{emory}$ model consistently outperforms the $\mathcal{N}$oise model across all metrics. Similarly, for Region-Set Captioning, both methods achieve strong results, but the $\mathcal{M}\textit{emory}$ method shows a clearer advantage, particularly in tasks closer to the patch level. Finally, in Image Captioning, the performance gap between the two architectures narrows, especially on the Flickr30k test split. In this scenario, the $\mathcal{M}\textit{emory}$ method performs significantly better when applied to the CLS token, whereas patch aggregation produces comparable results. However, the metrics reveal conflicting trends across different datasets.
@@ 13 Trace Captioning Benchmark Generation
We construct our Trace Captioning dataset from the Localized Narratives dataset [48]. This dataset consists of mouse traces and their corresponding speech transcriptions, where annotators describe objects in images while moving the mouse pointer over them.
@@
      {internallinenumbers*} I have image descriptions derived from spoken narratives. These need to be rewritten as concise, stand-alone captions in the style of the image-caption datasets. Follow these rules:      {internallinenumbers*} - Remove unnecessary narrative phrases like "we can see," "there is," "in this image," etc.    - Ensure the caption is standalone and descriptive.    - Use simple, objective language that highlights key elements.    - Keep it concise—just a single phrase.    - Follow the classical style of caption datasets.    {internallinenumbers*} - If the description is vague, subjective, or does not describe a concrete visual element (e.g., "The image is taken indoor," "This image is blurred"), return `<INVALID>`.    - Wrap the output in `{}` and add nothing else.      ### **Examples:**    - **Input:** "We can see a young elephant stands which is near the water in a wooded area."    **Output:** {A young elephant stands near the water in a wooded area.}      - **Input:** "In this image I can see some young children kicking a soccer ball in a field."    **Output:** {A group of young children kicking a soccer ball around a field.}      {internallinenumbers*} - **Input:** "In the left of the image, we see a pole that has two green street signs on it."    **Output:** {A pole has two green street signs on it.}      - **Input:** "We can see two surfboards which are stuck in the sand along the seashore."    **Output:** {Two surfboards stuck in the sand along the seashore.}      - **Input:** "This image consists of a man which rides a wakeboard behind a boat."    **Output:** {A man rides a wakeboard behind a boat.}      - **Input:** "In the background, there are a bunch of sticky notes and a pair of scissors."    **Output:** {A bunch of sticky notes and a pair of scissors.}      {internallinenumbers*} - **Input:** "It looks like a sepia-toned photograph of a motorcycle underneath the shadow of a    tree."    **Output:** {A sepia-toned photograph of a motorcycle underneath the shadow of a tree.}      - **Input:** "There is a sky"    **Output:** {A sky.}      - **Input:** "She is smiling."    **Output:** {A smiling girl.}      - **Input:** "The image is taken indoor."    **Output:** {<INVALID>}      - **Input:** "This image is edited."    **Output:** {<INVALID>}      - **Input:** "The image is blurred."    **Output:** {<INVALID>}      - **Input:** "I think he is about to jump."    **Output:** {<INVALID>}      Now, rewrite the following captions accordingly. Wrap each in `{}` and add nothing else:    <INPUT CAPTION>
@@
Figure 6: Narrative vs. Trace Samples. The first column displays sample images from the Localized Narrative dataset [49]. The remaining three columns show the corresponding mouse traces, along with the captions generated by the LLM. Captions marked with <INVALID> are removed from the dataset.
14 Learned Patch Aggregation via Attention
In the main paper we employ a parameter-free aggregation strategy to combine patch embeddings belonging to a region. While this choice preserves the zero-shot nature of our framework, we additionally explored whether a lightweight learned aggregation module could further improve the quality of region representations when local supervision is available.
Attention-based aggregation.
Let $S=\{v_{i}\}_{i=1}^{N}$, with $v_{i}\in\mathbb{R}^{D}$, denote the set of patch embeddings corresponding to a selected region. In the default formulation of our framework, the region representation is obtained through mean aggregation:
$v_{\text{mean}}=\frac{1}{N}\sum_{i=1}^{N}v_{i}.$ (3)
To investigate a learned alternative, we introduce a lightweight attention-based aggregation module that summarizes the patch set through a single transformer-style attention layer. The patch embeddings are processed jointly and produce a summary token that attends over the region patches. Formally, given the matrix of patch embeddings $V\in\mathbb{R}^{N\times D}$, we compute
$v_{\text{att}}=\mathrm{Attn}(V),$ (4)
where $\mathrm{Attn}(\cdot)$ denotes a self-attention block that outputs a single [CLS] token representing the aggregated region information.
Rather than replacing the mean representation, we combine the two signals to preserve the stability of the parameter-free aggregation while allowing the model to learn refinements. The final region embedding is defined as
$v_{R}=v_{\text{mean}}+\alpha\,v_{\text{att}},$ (5)
where $\alpha$ is a learnable scalar controlling the contribution of the attention-based summary. The parameter $\alpha$ is initialized to $0.1$, ensuring that the aggregation initially behaves close to the mean operator and gradually learns to incorporate the attention-based refinement during training.
Training setup.
The attention-based aggregator is trained with region-level supervision for a single epoch using the COCO Trace Captioning training set derived from the LocNar COCO train split. We optimize the parameters of the attention layer and the scalar weight $\alpha$ using AdamW with learning rate $10^{-4}$, while keeping the visual backbone and the text decoder frozen.
Model C P C P mAP C P mAP C P C P C P C P CLIP-S C P CLIP-S
Learned Aggregation (T2D + Noise) 73.9 81.9 34.8 77.2 17.8 49.4 81.7 17.7 48.9 81.7 67.8 83.6 25.7 73.3 74.2 86.5 70.9 27.3 79.7 66.0
Fixed Aggregation (T2D + Noise) 29.3 78.1 19.3 75.6 20.3 26.3 77.0 20.3 26.4 76.9 97.5 85.6 37.1 76.5 65.5 86.2 70.9 27.8 80.8 67.0
Table 11: Learned vs. Fixed Aggregation
Discussion.
Table 11 reports the results of this experiment. The learned aggregator yields substantial improvements when evaluated on tasks with the same spatial granularity used during training (trace and dense captioning). However, these gains do not consistently transfer to tasks involving different region granularities (e.g., region-set or image captioning), where the fixed aggregation remains more balanced.
These findings suggest that when task-specific local supervision is available, learned aggregation can effectively refine region representations. However, in a unified zero-shot framework where no region-level annotations are assumed, the parameter-free aggregation strategy remains preferable due to its robustness across captioning granularities.
@@ 15 More Qualitative Results
Additional qualitative results are shown in Figures 7 and 8. Note that the first rows of Figures 7 and 8 contain also qualitative results for single patch captioning, for which we do not have annotated data to report quantitative results.
As can be noticed in Figures 7 and 8, the Region-Set Captioning task tends to align more closely with image-level captioning rather than strictly focusing on localized regions. This is expected since the ground-truth captions in the COCO Entities dataset originate from the image-level annotations of COCO, as stated in [10].
@@
Figure 7: Qualitative results. We report four predictions of our model and compare baselines from the finer (top) to the coarser (bottom) task. For trace captioning examples, the trace time is color-coded from start (red) to end (yellow). DeCap = DeCap applied on the whole image. DeCap (Crop) = DeCap applied on cropped box. ZeroCap = ZeroCap [59] applied to the whole image. CLOSE = CLOSE [18] applied to the whole image. Ours (CLIP + Mem.) = Our patch-based framework using CLIP as backbone and the projection as modality gap mitigation strategy. Ours (Talk2DINO + Mem.) = Our patch-based framework using Talk2DINO as backbone and the projection as modality gap mitigation strategy. GT = ground-truth caption.
@@
Figure 8: Qualitative results. We report four predictions of our model and compare baselines from the finer (top) to the coarser (bottom) task. For trace captioning examples, the trace time is color-coded from start (red) to end (yellow). DeCap = DeCap applied on the whole image. DeCap (Crop) = DeCap applied on cropped box. ZeroCap = ZeroCap [59] applied to the whole image. CLOSE = CLOSE [18] applied to the whole image. Ours (CLIP + Mem.) = Our patch-based framework using CLIP as backbone and the projection as modality gap mitigation strategy. Ours (Talk2DINO + Mem.) = Our patch-based framework using Talk2DINO as backbone and the projection as modality gap mitigation strategy. GT = ground-truth caption.